### v1

# DMS*: Minimizing Makespan for Multi-Agent Combinatorial Path Finding

Abstract
Multi-Agent Combinatorial Path Finding (MCPF) seeks collision-free paths for multiple agents from their initial to goal locations, while visiting a set of intermediate target locations in the middle of the paths. MCPF is challenging as it involves both planning collision-free paths for multiple agents and target sequencing, i.e., solving traveling salesman problems to assign targets to and find the visiting order for the agents. Recent work develops methods to address MCPF while minimizing the sum of individual arrival times at goals. Such a problem formulation may result in paths with different arrival times and lead to a long makespan, the maximum arrival time, among the agents. This paper proposes a min-max variant of MCPF, denoted as MCPF-max, that minimizes the makespan of the agents. While the existing methods (such as MS*) for MCPF can be adapted to solve MCPF-max, we further develop two new techniques based on MS* to defer the expensive target sequencing during planning to expedite the overall computation. We analyze the properties of the resulting algorithm Deferred MS* (DMS*), and test DMS* with up to 20 agents and 80 targets. We demonstrate the use of DMS* on differential-drive robots.
Index Terms:
Path Planning for Multiple Mobile Robots or Agents, Multi-Agent Path Finding, Traveling Salesman Problem
I Introduction
Multi-Agent Path Finding (MAPF) seeks a set of collision-free paths for multiple agents from their respective start to goal locations. This paper considers a generalization of MAPF called Multi-Agent Combinatorial Path Finding (MCPF), where the agents need to visit a pre-specified set of intermediate target locations before reaching their goals. MAPF and MCPF arise in applications such as surveillance [1] and logistics [2]. For instance, factories [3, 4] use a fleet of mobile robots to visit a set of target locations to load machines for manufacturing. These robots share a cluttered environment and follow collision-free paths. In such settings, MAPF problems and their generalizations naturally arise to optimize operations.
MCPF is challenging as it involves both collision avoidance among the agents as present in MAPF, and target sequencing, i.e., solving Traveling Salesman Problems (TSPs) [5, 6] to specify the allocation and visiting orders of targets for all agents. Both the TSP and the MAPF are NP-hard to solve to optimality [7, 5], and so is MCPF. A few methods have been developed [8, 9] to address the challenges in MCPF, and they often formulate the problem as a min-sum optimization problem, denoted as MCPF-sum, where the objective is to minimize the sum of individual arrival times. Such a formulation may result in an ensemble of paths where some agents arrive early while others arrive late, which leads to long execution times before all agents finish their paths. This paper thus proposes a min-max variant of MCPF (Fig. 1), denoted as MCPF-max, where the objective is to minimize the maximum arrival time, which is also called the makespan, of all agents.
To solve MCPF-max, this paper first adapts our prior MS* algorithm [8], which was first designed for MCPF-sum, to address MCPF-max. Then, we further develop two new techniques to expedite the planning, and we call the resulting new algorithm Deferred MS* (DMS*).
Figure 1: MCPF-max problem, which seeks a set of collision-free paths while minimizing the maximum arrival time of all agents. The color of a target or goal indicates the assignment constraints, i.e., the subset of agents that are eligible to visit that target or goal. The orange target appears on the path of the blue agent, but is not claimed to be visited by the blue agent.
Specifically, the existing MS* follows the general heuristic search framework (such as A*) by iteratively generating, selecting and expanding states to construct partial solution paths from the initial state to the goal state. MS* uses Traveling Salesman Problem (TSP) algorithms to compute target sequences for agents. When solving the TSP, agent-agent collision are ignored, and the cost of resulting target sequences are thus lower bounds of the true costs for the agents to arrive at their goals. Therefore, the cost of target sequences provides an admissible heuristic to guide the state selection and expansion as in A*. Furthermore, MS* leverages the idea in M* [10] to first use the target sequences to build a low-dimensional search space, and then grow this search space by coupling agents together for planning only when collision happens. By doing so, MS* interleaves TSP (target sequencing) and MAPF (collision resolution) techniques under the heuristic search framework and is able to provide completeness and solution quality guarantees.
The first technique developed in this paper is applicable to MS* for both MCPF-max and MCPF-sum. When expanding a state, a set of successor states are generated, and for each of them, MS* needs to invoke the TSP solver to find the target sequence and the heuristic value of this successor state. Since the number of successor states can be large for each expansion, MS* needs to frequently invokes the TSP solver which slows down the computation. To remedy this issue, for each generated successor, we first use a fast-to-compute yet roughly estimated cost-to-go as the heuristic value, and defer calling TSP solver for target sequencing until that successor is selected for expansion.
The second technique developed in this paper is only applicable to MS* when solving MCPF-max, and does not work for MCPF-sum. Since the goal here is to minimize the makespan, during the search, agents with non-maximum arrival time naturally have “margins” in a sense that they can take a longer path without worsening the makespan of all agents. We take advantage of these margins to let agents re-use their previously computed target sequences and defer the expensive calls of TSP solvers until the margin depletes.
To verify the methods, we conduct both simulation in various maps with up to 20 agents and 80 targets, as well as real robot experiments. The simulation shows that the new techniques in DMS* help triple the success rates and reduce the average runtime to solution. The real robot experiments demonstrate that the planned path are executable on real robots, and that DMS* is able to optimize makespan while the existing planners that minimize the sum of arrival times can lead to long makespan.
II Related Work
II-A Multi-Agent Path Finding
To find collision-free paths for multiple agents, a variety of MAPF algorithms were developed, falling on a spectrum from coupled [11] to decoupled [12], trading off completeness and optimality for scalability. In the middle of this spectrum lie the dynamically-coupled methods such as M* [10] and CBS [13], which begin by planning for each agent a shortest path from the start to the goal ignoring any potential collision with the other agents, and then couple agents for planning only when necessary in order to resolve agent-agent collision. These dynamically-coupled methods were improved and extended in many ways [14, 15, 16, 17].
II-B Traveling Salesman Problems
In the presence of multiple intermediate targets, the planner needs to determine both the assignment and visiting order of the targets for the agents. For a single agent, the well-known Traveling Salesman Problem (TSP), which seeks a shortest tour that visits every vertex in a graph, is one of the most well-known NP-hard problems [5]. Closely related to TSP, the Hamiltonian Path Problem (HPP) requires finding a shortest path that visits each vertex in the graph from a start vertex to a goal vertex. The multi-agent version of the TSP and HPP (denoted as mTSP and mHPP, respectively) are more challenging since the vertices in the graph must be allocated to each agent in addition to finding the optimal visiting order of vertices. To simplify the presentation, we refer to all these problems simply as TSPs. To solve TSPs, a variety of methods have been developed ranging from exact methods (branch and bound, branch and price) [5] to heuristics [18] and approximation algorithms [19], trading off solution optimality for runtime efficiency. This work does not develop new TSPs solvers and leverage the existing ones.
II-C Target Assignment, Sequencing and Path Finding
Some recent research combines MAPF with target assignment and ordering [20, 21, 22, 23, 24, 25, 9, 8, 26, 27, 28, 29]. Most of them either consider target assignment only (without the need for computing visiting orders of targets) [20, 21, 22], or consider the visiting order only given that each agent is pre-allocated a set of targets [23, 24, 25]. Our prior work [8, 9] seeks to handle the challenge in target assignment and ordering as well as the collision avoidance in MAPF simultaneously. These work uses the MCPF-sum formulation and the developed planners minimize the sum of individual arrival times, while this paper investigates MCPF-max, which seeks to minimize the maximum arrival time.
III Problem
Let index set $I=\{1,2,\dots,N\}$ denote a set of $N$ agents. All agents share a workspace that is represented as an undirected graph $G^{W}=(V^{W},E^{W},c^{W})$, where $W$ stands for workspace. Each vertex $v\in V^{W}$ represents a possible location of an agent. Each edge $e=(u,v)\in E^{W}\subseteq V^{W}\times V^{W}$ represents an action that moves an agent between $u$ and $v$. $c^{W}:E^{W}\rightarrow(0,\infty)$ maps an edge to its positive cost value. In this paper, the cost of an edge is equal to its traversal time, and each edge has a unit cost. [^1] In [8, 9], the edge cost can be different from its traversal time: the edge cost can be any positive number while the traversal time is always one unit per edge. In this paper, the edge cost are the same as the edge traversal time, which is one for each edge.
Let the superscript $i\in I$ over a variable denote the specific agent to which the variable belongs (e.g. $v^{i}\in V^{W}$ means a vertex corresponding to agent $i$). Let $v_{o}^{i},v_{d}^{i}\in V^{W}$ denote the initial (or original) vertex and the goal (or destination) vertex of agent $i$ respectively. Let $V_{o},V_{d}\subset V^{W}$ denote the set of all initial and goal vertices of the agents respectively, and let $V_{t}\subset V^{W}$ denote the set of target vertices. For each target $v\in V_{t}$, let $f_{A}(v)\subseteq I$ denote the subset of agents that are eligible to visit $v$; these sets are used to formulate the (agent-target) assignment constraints.
Let $\pi^{i}(v^{i}_{1},v^{i}_{\ell})$ denote a path for agent $i$ between vertices $v^{i}_{1}$ and $v^{i}_{\ell}$, which is a list of vertices $(v^{i}_{1},v^{i}_{{2}},\dots,v^{i}_{\ell})$ in $G^{W}$ with $(v^{i}_{k},v^{i}_{k+1})\in E^{W},k=1,2,\cdots,\ell-1$. Let $g(\pi^{i}(v^{i}_{1},v^{i}_{\ell}))$ denote the cost of the path, which is the sum of the costs of all edges present in the path: $g(\pi^{i}(v^{i}_{1},v^{i}_{\ell}))=\Sigma_{j=1,2,\dots,{\ell-1}}c^{W}(v^{i}_{{j}},v^{i}_{{j+1}})$.
All agents share a global clock and start to move along their paths from time $t=0$. Each action of the agents, either wait or move along an edge, requires one unit of time. Any two agents $i,j\in I$ are in conflict if one of the following two cases happens. The first case is a vertex conflict $(i,j,v,t)$ where two agents $i,j\in I$ occupy the same vertex $v$ at the same time $t$. The second case is an edge conflict $(i,j,e,t)$, where two agents $i,j\in I$ go through the same edge $e$ from opposite directions between times $t$ and $t+1$.
Definition 1 (MCPF-max Problem).
The Multi-Agent Combinatorial Path Finding with Min-Max Objective (MCPF-max) seeks to find a set of conflict-free paths for the agents such that (1) each target $v\in V_{t}$ is visited [^2] In MCPF-max, the notion that an agent $i$ “visits” a target $v\in V_{t}$ means (i) there exists a time $t$ such that agent $i$ occupies $v$ along its path, and (ii) the agent $i$ claims that $v$ is visited. In other words, if a target $v$ is in the middle of the path of agent $i$ and agent $i$ does not claim $v$ is visited, then $v$ is not considered as visited. Additionally, a visited target $v$ can appear in the path of another agent. For the rest of the paper, when we say an agent or a path “visits” a target, we always mean the agent “visits and claims” the target. at least once by some agent in $f_{A}(v)$, (2) the path for each agent $i\in I$ starts at its initial vertex and terminates at a unique goal vertex $v^{i}_{d}\in V_{d}$ such that $i\in f_{A}(v^{i}_{d})$, and (3) the maximum of the cost of all agents’ paths (i.e., $\max_{i\in I}cost(\pi^{i}(v^{i}_{o},v^{i}_{d})))$ reaches the minimum.
Remark 1.
The MCPF-max is similar to MCPF-sum [9, 8]. The only difference between these two problems is that: MCPF-sum seeks to minimize the sum of individual path costs while MCPF-max minimizes the maximum of individual path costs.
IV Method
This section begins with a toy example in Fig. 2 to illustrate the planning process of DMS*, and provides the explanation in the caption. This section then introduces the concepts and notations, explains the pseudo-code and elaborates the technical detail.
Figure 2: An illustration of DMS* and related concepts. (a) shows the workspace graph $G^{W}$ and a toy problem. (b) shows the corresponding target graph $G^{T}$ where each edge in $G^{T}$ corresponds to a minimum cost path in $G^{W}$ between the respective vertices. (c) shows the workflow of DMS* in Alg. 1. (d) shows that DMS* first ignores any agent-agent conflict and solves a corresponding mHPP, which provides a joint sequence $\gamma(l_{o})$ (d1). This joint sequence $\gamma(l_{o})$ can be converted to a joint path $\pi(l_{o})$ (d2,d3), whose corresponding makespan $f_{max}$ is 9. (e) The joint path $\pi(l_{o})$ leads to a policy $\phi_{\gamma(l_{o})}$ that maps a joint vertex to the next joint vertex along $\pi(l_{o})$. Since agent-agent conflicts are ignored, when following the policy, agents may run into conflict (e2). When a conflict is detected, the subset of agents that are in conflict $I_{C}(l)$ is back propagated to the nearby ancestor labels. (f) After the back propagation, these ancestor labels and re-opened and re-expanded while considering all possible actions of the agents that are in conflict (f1,f2,f3). During this re-expansion, for each of the successor, DMS* first uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic of the generated labels to avoid solving a mHPP. For both labels (f2) and (f3), this rough heuristic is $(9,5)-(1,1)=(8,4)$, where $(9,5)$ is the heuristic of (f1), the parent of (f2) and (f3), and $(1,1)$ means that each agent can move at most one step towards their goals. When either of (f2) and (f3) is popped from OPEN and before being expanded, DMS* re-computes a new heuristic and see if the popped label should be expanded or re-added to OPEN for future expansion. To obtain this new heuristic, DMS* first attempts to let the agents follow the previously computed target sequences of the parent label and check if this will worsen the overall makespan. For the successor shown in (f2), the resulting makespan is 9 (f4), which is no worse than the previous makespan 9, and DMS* will not call mHPP solver to save computational effort. For the successor shown in (f3), the resulting makespan is 10 (f5), which is worse than the previous makespan 9, and DMS* have to call mHPP solver to find an optimal joint sequence from that successor (f3) to the goals. Finally, with the joint sequence, new policy for the successors can be built and the search continues as in (e).
IV-A Concepts and Notations
IV-A1 Joint Graph
Let $\mathcal{G^{W}}=(\mathcal{V^{W}},\mathcal{E^{W}})=\underbrace{G^{W}\times G^{W}\times\dots\times G^{W}}_{\text{$N$ times}}$ denote the joint graph of all agents which is the Cartesian product of $N$ copies of $G^{W}$, where each vertex $v\in\mathcal{V^{W}}$ represents a joint vertex and $e\in\mathcal{E^{W}}$ represents a joint edge that connects a pair of joint vertices. The joint vertex $v_{o}=(v^{1}_{o},v^{2}_{o},\cdots,v^{N}_{o})$ corresponding to the initial vertices of all agents is called the initial joint vertex. Let $\pi(u,v),u,v\in\mathcal{V^{W}}$ denote a joint path, which is a tuple of $N$ (individual) paths of the same length, i.e., $\pi(u,v)=(\pi^{1}(u^{1},v^{1}),\cdots,\pi^{N}(u^{N},v^{N}))$. For the rest of this paper, we use a “path” to denote an “individual path”. DMS* searches the joint graph $\mathcal{G^{W}}$ for a joint path that solves the MCPF-max.
IV-A2 Binary Vector
For any joint vertex $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g. $\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited along the joint path, and we need to differentiate between them during the search. We therefore introduce the following concepts. First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{t_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list. [^3] In this paper, for a vector related to targets (e.g. a binary vector $\vec{a}$ of length $M$), we use subscripts (e.g. $a_{m}$) to indicate the elements in the vector. For a vector or joint vertex that is related to agents (e.g. $\vec{g}$ of length $N$, or $v\in\mathcal{V^{W}}$), we use a superscript (e.g. $i$ in $g^{i},v^{i}$) to indicate the element in the vector or joint vertex corresponding to an agent. Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if the corresponding target $t_{m}$ is visited by some agent, and $a_{m}=0$ otherwise.
IV-A3 Label
Let $l=(v,\vec{a},\vec{g})$ denote a label, where $v$ is a joint vertex, $\vec{a}$ is a binary vector of length $M$, and $\vec{g}$ is a cost vector of length $N$. Here, each component $g^{i},i\in I$ corresponds to the path cost of agent $i$. Intuitively, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path costs as specified by $\vec{g}$. To simplify the presentation, given a label $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$. Furthermore, let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$. To solve the MCPF-max in Def. 1, the algorithm needs to find a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reach the goals, and $g_{max}(l)$ is the objective value to be minimized.
IV-A4 Label Comparison
To compare two labels at the same joint vertex (i.e., to compare the joint paths represented by these two labels), we need to compare both their binary vectors as well as the path costs. We introduce the following dominance relationship to compare two binary vectors.
Definition 2 (Binary Dominance).
For any two binary vectors $\vec{a}$ and $\vec{b}$, $\vec{a}$ dominates $\vec{b}$ ($\vec{a}\succeq_{b}\vec{b}$), if both the following conditions hold: (i) $\forall m\in\{1,2,\cdots,M\}$, ${a}_{m}\geq{b}_{m}$; (ii) $\exists m\in\{1,2,\cdots,M\}$, ${a}_{m}>{b}_{m}$.
Intuitively, $\vec{a}$ dominates $\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$. In addition, two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
To compare two labels with the same joint vertex, we define the following dominance rule between them.
Definition 3 (Label Dominance).
For any two labels ${l}_{1}$ and ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either of the following two conditions holds: (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$.
Intuitively, if $l_{1}$ dominates $l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$. If $l_{1}$ does not dominate $l_{2}$, $l_{1}$ is then non-dominated by $l_{2}$. Any two labels are non-dominated (with respect to each other) if each of them is non-dominated by the other. Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if they have the same vertex, binary vector and $g_{max}$ (i.e., $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$). Note that there is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Def. III seeks to minimize the maximum path cost over the agents. Finally, for each joint vertex $v\in\mathcal{V^{W}}$, let $\mathcal{F}(v)$ denote a set of labels that are non-dominated to each other during the search.
IV-A5 Target Sequencing
Let $\gamma^{i}=\{v^{i}_{o},t_{1},t_{2},\cdots,t_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where $v^{i}_{o}\in V_{o}$ is the initial vertex of agent $i\in I$, each $t_{j}\in V_{t},j=1,2,\cdots,k$ is a target vertex and $v^{i}_{d}\in V_{d}$ is a goal vertex. A target sequence $\gamma^{i}$ specifies both the assignment and visiting order of targets for agent $i$. Let $\gamma=\{\gamma^{1},\gamma^{2},\cdots,\gamma^{N}\}$ denote a joint sequence, which specify the assignment and visiting order of all targets for all agents.
Given a label $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited. We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, which is a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$. In other words, (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$, the vertex of agent $i$ in $v(l)$, visits a set of unvisited targets $\{t_{m}\}$, where the $m$-th component of $\vec{a}(l)$ is zero, and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ collectively visits all unvisited targets as specified in $\vec{a}(l)$.
IV-A6 Target Graph
Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in the workspace graph $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$. The cost of a target sequence $cost(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$.
Given a label $l$, to find a joint sequence $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved. Let SolveMHPP($l$) denote a procedure that solves the mHPP. SolveMHPP($l$) first formulates a target graph $G^{T}=(V^{T},E^{T},c^{T})$, which is computed based on the workspace graph $G^{W}$ as follows. The vertex set $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{t_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$. $G^{T}$ is fully connected and $E^{T}=V^{T}\times V^{T}$. The edge cost in $G^{T}$ of any pair of vertices $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the workspace graph $G^{W}$. An example $G^{T}$ is shown in Fig. 2(b). The aforementioned target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$ and a joint sequence $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as specified by $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints. The procedure SolveMHPP$(l)$ can be implemented by various existing algorithms for mHPP.
IV-A7 Heuristic and Policy
Given a joint sequence $\gamma(l)$, to simplify the notations, let $h^{i}(l):=cost(\gamma^{i}(l))$ be a heuristic value. Since agent-agent conflicts are ignored along the target sequences, the vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$. When SolveMHPP($l$) solves the mHPP to optimality, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which can be used as an admissible heuristic for the search. For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from the initial vertices to goals for all agents by further extending the joint path represented by $l$. Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to label $l$.
Given a label $l$ and its joint sequence $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ can be built out of $\gamma(l)$, which maps one label to another label as follows. First, a joint path $\pi$ can be built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l)$ for any agent $i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$. Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from one joint vertex $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$. A set of corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ can be built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by considering if $v_{k}$ visits any new targets. A set of corresponding cost vectors $\vec{g}_{k}$ can also be computed in a similar way as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$. As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l_{k}=(v_{k},\vec{a}_{k},\vec{g}_{k})$ to the next label $l_{k+1}=(v_{k+1},\vec{a}_{k+1},\vec{g}_{k+1})$ along the target sequence $\phi_{\gamma(l)}$. Given a vertex $v^{i}$ of agent $i$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$. An example of $\phi^{i}_{\gamma(l)}$ is shown in Fig. 2(d). A label $l$ is called on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$ for some $l^{\prime}$. Otherwise, $l$ is called off-policy, which means its next label is unknown yet and a mHPP needs to be solved from $l$ to find the policy $\phi_{\gamma(l)}$. We denote $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) as the process of computing the joint sequence, heuristic values and joint policy.
IV-B DMS* Algorithm
Algorithm 1 Pseudocode for DMS*
[…]
2:$parent(l_{o})\leftarrow NULL$, $f_{temp}(l_{o})\leftarrow 0$
3:add $l_{o}$ to OPEN with $f_{temp}(l_{o})$ as the priority
4:add $l_{o}$ to $\mathcal{F}(v_{o})$
5:while OPEN is not empty do
6:    $l=(v,\vec{a},\vec{g})\leftarrow$ OPEN.pop()
7:    TargetSeq($parent(l)$, $l$, $I_{C}(parent(l))$)
8:    $f_{max}(l)\leftarrow\max_{i\in I}\{\vec{g}(l)+w\cdot\vec{h}(l)$}
9:    if $f_{max}(l)>f_{temp}(l)$then
10:         $f_{temp}(l)\leftarrow f_{max}(l)$
11:         add $l$ to OPEN with $f_{temp}(l)$ as the priority
12:         continue
[…]
14:    return Reconstruct($l$)
15:    $L_{succ}\leftarrow$ GetSuccessors($l$)
16:    for all $l^{\prime}\in L_{succ}$ do
[…]
18:         BackProp($l$, $I_{C}(l^{\prime})$)
19:         if $I_{C}\neq\emptyset$ then
20:             continue
21:         if IsDominated($l^{\prime}$then
22:             DomBackProp($l,l^{\prime}$)
23:             continue
[…]
25:         $parent(l^{\prime})\leftarrow l$
26:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
27:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
28:return Failure (no solution)
Algorithm 2 Pseudocode for TargetSeq($l$, $l^{\prime}$, $I_{C}(l)$)
1:INPUT: two labels $l,l^{\prime}$ where $l$ is the parent of $l^{\prime}$.
2:if $l^{\prime}$ is on-policy then
3:    $\gamma(l^{\prime})\leftarrow\gamma(l)$
4:    return
5:$\gamma(l^{\prime})\leftarrow\gamma(l)$
6:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
7:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
[…]
9:return
Algorithm 3 Pseudocode for BackProp($l,I_{C}(l^{\prime})$)
1:if $I_{C}(l^{\prime})\nsubseteq I_{C}(l)$ then
2:    $I_{C}(l)\leftarrow I_{C}(l^{\prime})\bigcup I_{C}(l)$
3:    if $l\notin$ OPEN then
4:         add $l$ to OPEN
5:    for all $l^{\prime\prime}\in$ back_set($l$do
6:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
Algorithm 4 Pseudocode for DomBackProp($l,l^{\prime}$)
1:for all $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ do
2:    if $l^{\prime\prime}\succeq_{l}l^{\prime}\text{ or }l^{\prime\prime}=l^{\prime}$ then
3:         BackProp($l$, $I_{C}(l^{\prime\prime})$)
4:         add $l$ to back_set($l^{\prime\prime}$)
To initialize (Lines 1-3), DMS* begins by creating an initial label $l_{o}$ and invokes the procedure SolveMHPP (as aforementioned) to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$. For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$. Here, $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals. In contrast, $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$. Similarly to A* [30], let OPEN denote a priority queue that stores labels $l$ and prioritizes them based on their $f$-values from the minimum to the maximum. In Alg. 1, we point out which $f$-value (either $f_{temp}$ or $f_{max}$) is used when a label is added to OPEN. Finally, $l_{o}$ is added to $\mathcal{F}(v_{o})$ since $l_{o}$ is non-dominated by any other labels at $v_{o}$, and $l_{o}$ is added to OPEN for future search. [^4] The notion of frontier set in this paper differs from the ones in . In , $\mathcal{F}(v)$ contains only expanded labels, which resembles the closed set in conventional A*, while in this paper, $\mathcal{F}(v)v\in\mathcal{V^{W}}$ includes both the labels that are expanded (closed labels) and labels that are generated and not yet expanded (open labels).
In each search iteration (Lines 5-27), DMS* begins by popping a label $l$ from OPEN for processing. DMS* first calls the TargetSeq procedure for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$. TargetSeq determines either to call SolveMHPP to find a joint sequence for $l$ (i.e., $\gamma(l)$), or to re-use the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$ as $\gamma(l)$. We elaborate TargetSeq in Sec. IV-C. After TargetSeq on Line 7, the heuristic $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes the $f_{max}(l)$ and compare it against $f_{temp}(l)$. When computing $f_{max}$ out of $\vec{g}$ and $\vec{h}$ on Line 8, a heuristic inflation factor $w\in[1,\infty)$ is used, which scales each component in $\vec{h}$ by the factor $w$. Heuristic inflation is a common technique for A* [31] and M*-based algorithms [10] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [31]. If $f_{max}(l)>f_{temp}(l)$, then $l$ should not be expanded in the current iteration since there can be labels in OPEN that have smaller $f$-value than $f_{max}(l)$. DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and skip the current search iteration. In a future search iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
Afterwards, DMS* checks if $l$ leads to a solution using CheckSuccess($l$), which verifies if every component in $\vec{a}(l)$ is one and if every component of $v(l)$ is a unique goal vertex while satisfying the assignment constraints. If CheckSuccess($l$) returns true, a solution joint path $\pi^{*}$ is found and can be reconstructed by iterative tracking the parent pointers of labels from $l$ to $l_{o}$ in Reconstruct($l$). DMS* then terminates.
If CheckSuccess($l$) returns false, $l$ is expanded by considering its limited neighbors [10] described as follows. Let $I_{C}(l)\subseteq I$ denote the conflict set of label $l$, which is a subset of agents that are (or will be) in conflict. The limited neighbors of $l$ is a set of successor labels of $l$. For each agent $i$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$. If $i\in I_{C}(l)$, agent $i$ is allowed to visit any adjacent vertex of $v^{i}(l)$ in $G^{W}$. In other words, the successor vertices of $v^{i}(l)$ are the following:
$u^{i}\leftarrow\begin{cases}\phi^{i}_{\gamma(l)}(v^{i}(l))&\mbox{if }i\notin I_{C}(l)\\ u^{i}\;|\;(v^{i}(l),u^{i})\in E^{W}&\mbox{if }i\in I_{C}(l)\end{cases}$ (1)
Let $V_{succ}^{i}$ denote the set of successor vertices of $v^{i}(l)$, which is either of size one or equal to the number of edges incident on $v^{i}(l)$ in $G^{W}$. The successor joint vertices $V_{succ}$ of $v(l)$ is then the combination of $v^{i}(l)$ for all $i\in I$, i.e., $V_{succ}:=V_{succ}^{1}\times V_{succ}^{2}\times\cdots\times V_{succ}^{N}$. For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding label $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$. When creating $l^{\prime}$, the corresponding $\vec{g}(l^{\prime})$ and $\vec{a}(l^{\prime})$ are computed based on $\vec{g}(l)$ $\vec{a}(l)$ and $v^{\prime}$.
After generating the successor labels $L_{succ}$ of $l$, for each label $l^{\prime}\in L_{succ}$, DMS* checks for conflicts between agents during the transition from $v(l)$ to $v(l^{\prime})$, and store the subset of agents in conflict in the conflict set $I_{C}(l^{\prime})$. DMS* then invokes a procedure BackProp (Alg. 3) to back propagate $I_{C}(l^{\prime})$ to its ancestor labels recursively so that the conflict set of these ancestor labels are modified, and labels with modified conflict set are re-added to OPEN and will be re-expanded. DMS* maintains a back_set($l$) for each label $l$, which is a set of pointers pointing to the predecessor labels to which the back propagation should be conducted. Intuitively, similarly to [10, 8], the conflict set of labels are dynamically enlarged during planning when agents are detected in conflict. The conflict sets of labels determine the sub-graph within the joint graph $\mathcal{G^{W}}$ that can be reached by DMS*, and DMS* always attempts to limit the search within a sub-graph of $\mathcal{G^{W}}$ as small as possible.
Afterwards, if the conflict set $I_{C}(l^{\prime})$ is non-empty, then label $l^{\prime}$ leads to a conflict and is discarded. Otherwise, $l^{\prime}$ is checked for dominance against any existing labels in $\mathcal{F}(v(l^{\prime}))$ using Def. 3. If $l^{\prime}$ is dominated by or is equal to any existing labels in $\mathcal{F}(v(l^{\prime}))$, $l^{\prime}$ is pruned, since any future joint path from $l^{\prime}$ can be cut and paste to a label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates $l^{\prime}$ without worsening the cost to reach goals. Furthermore, for each label $l^{\prime\prime}\in\mathcal{F}(v(l^{\prime}))$ that dominates or is equal to $l^{\prime}$, the procedure DomBackProp(Alg. 4) is invoked so that the conflict set $I_{C}(l^{\prime})$ is back propagated to $l$, and $l$ is added to the back_set of $l^{\prime\prime}$. By doing so, DMS* is able to keep updating the conflict set of the predecessor labels of $l^{\prime}$ after $l^{\prime}$ is pruned. This ensures that, when needed, the predecessor labels of $l^{\prime}$ will also be re-expanded after $l^{\prime}$ is pruned.
If label $l^{\prime}$ is not pruned, a procedure SimpleHeu is invoked for $l^{\prime}$ to quickly compute a heuristic that roughly estimates the cost-to-go. One possible way to implement SimpleHeu is to first copy $\vec{h}(l)$, where $l$ is the parent of $l^{\prime}$, and then reduce each component of the copied vector by one except for the components that are already zero. This heuristic is still an underestimate of the cost-to-go since all agents can move at most one step closer to their goals in each expansion. Then, the $f_{temp}(l^{\prime})$ can be computed and $l^{\prime}$ is added to OPEN with $f_{temp}(l^{\prime})$ as its priority. Other related data structure including $\mathcal{F}(v(l^{\prime}))$, back_set, $parent$ are also updated correspondingly, and the current search iteration ends.
When DMS* terminates, it either finds a conflict-free joint path, or returns failure when OPEN is empty if the given instance is unsolvable.
IV-C Deferred Target Sequencing
DMS* introduces two techniques to defer the target sequencing until absolutely needed. As aforementioned, the first one uses a fast-to-compute yet roughly estimated cost-to-go as the heuristic when a label is generated, and invokes TargetSeq only when that label is popped from OPEN for expansion.
We now focus on the second technique in DMS*. In the previous MS* [8], every time when the search encounters a new label $l$ that is off-policy, inside TargetSeq, the procedure SolveMHPP needs to be invoked for $l$ to find a joint sequence and policy from $l$, which makes the overall computation burdensome, especially when a lot of new labels are generated due to the agent-agent conflict.
Different from MS*, DMS* seeks to defer the call of SolveMHPP inside TargetSeq. For a label $l^{\prime}$, DMS* attempts to avoid calling SolveMHPP for $l^{\prime}$ by re-using the joint sequence $\gamma(l)$ of its parent label $l$ ($l$ is the parent of $l^{\prime}$). As presented in Alg. 2, on Lines 5-6, DMS* first attemps to build a policy from $l^{\prime}$ by following the joint sequence of its parent label $\gamma(l)$ and computes the corresponding cost-to-go $\vec{h}(l^{\prime})$. Agents $i$ that are not in the conflict set $I_{C}(l)$ (i.e., $i\notin I_{C}(l)$) are still along their individual paths as specified by the previously computed policy $\phi_{\gamma(l)}$. Agents $i$ that are in the conflict set $I_{C}(l)$ (i.e., $i\in I_{C}(l)$) consider all possible actions as described in Equation (1) and may deviate from the individual paths specified by the previously computed policy $\phi_{\gamma(l)}$. Therefore, DMS* needs to go through a check for these agents $i\in I_{C}(l)$: DMS* first computes the $f$-vector by summing up $\vec{g}(l^{\prime})$ and $\vec{h}(l^{\prime})$. Then, if $f^{i}(l^{\prime})$ of some agent $i\in I_{C}(l)$ is no larger than $f_{max}(l)$, DMS* can avoid calling SolveMHPP, since letting the agents follow $\gamma(l)$ in the future will not worsen the objective value. Otherwise, there exists an agent $i\in I_{C}(l)$, whose corresponding $f^{i}(l^{\prime})$ is greater than $f_{max}(l)$, and in this case, DMS* cannot avoid calling SolveMHPP since there may exists another joint sequence from $l^{\prime}$, which leads to a solution joint path with better (smaller) objective value.
IV-D Properties of DMS*
IV-D1 Completeness
This section discusses the properties of DMS*. An algorithm is complete for a problem if finds a solution for solvable instances, and reports failure in finite time for unsolvable instances.
Theorem 1.
DMS* is complete for MCPF-max.
Let $\mathcal{A}=\{0,1\}^{M}$ denote the set of all possible binary vectors and let $\mathcal{S}=\mathcal{G^{W}}\times\mathcal{A}$ denote the Cartesian of the joint graph $\mathcal{G^{W}}$ and $\mathcal{A}$. $\mathcal{S}$ is a finite space. DMS* uses dominance pruning and does not expand the same label twice during the search. As a result, for a unsolvable instance, DMS* enumerates all possible joint paths that starts from $(v_{o},\vec{0}^{M})$ to any reachable $(v,a)\in\mathcal{S}$ and terminates in finite time. For a solvable instance, DMS* conducts systematic search in $\mathcal{S}$ and returns a solution in finite time.
IV-D2 Solution Optimality
To discuss the solution optimality of DMS*, we needs the following assumptions.
Assumption 1.
The procedure SolveMHPP returns an optimal joint sequence for the given mHPP instance.
Assumption 2.
Line 7 in Alg. 1 never returned true during the search of DMS*.
Theorem 2.
For a solvable instance, when Assumption 1 and 2 hold, DMS* returns an optimal solution joint path.
The proof of Theorem 2 follows the analysis in [10, 8], and we highlight the main ideas here. The policies computed by DMS* defines a sub-graph $\mathcal{G}^{W}_{sub}$ of the joint graph $\mathcal{G^{W}}$. DMS* first expands labels in this sub-graph $\mathcal{G}^{W}_{sub}$. If no conflict is detected when following the policy, then the resulting joint path is conflict-free and optimal due to Assumption 1: With Assumption 1, ${h}_{max}(l)$ for any label $l$ computed by SolveMHPP is an estimated cost-to-go that is admissible, i.e., ${h}_{max}(l)$ is no larger than the true optimal cost-to-go. This is true because SolveMHPP ignores agent-agent conflicts and the resulting ${h}_{max}(l)$ must be a lower bound on the true cost-to-go. As DMS* selects labels from OPEN in the same way as A* does, admissible heuristics lead to an optimal solution [31].
If conflicts are detected, DMS* updates (i.e., enlarges) $\mathcal{G}^{W}_{sub}$ by enlarging the conflict set and back propagating the conflict set. When Assumption 2 holds, the enlarged $\mathcal{G}^{W}_{sub}$ still ensures that an optimal conflict-free joint path $\pi_{*}$ is contained in the updated $\mathcal{G}^{W}_{sub}$ [10, 8]. DMS* systematically search over $\mathcal{G}^{W}_{sub}$ and finds $\pi_{*}$ at termination.
We now explain if Assumption 2 is violated, why DMS* loses the optimality guarantee and how to fix it. In MCPF-max, all agents are “coupled” in the space of binary vectors $\mathcal{A}$ in a sense that a target visited by one agent does not need to be visited by any other agents. As as result, when one agent changes its target sequence by visiting some target that is previously assigned to another agent, all agents may need to be re-planned in order to ensure solution optimality. It means that, when Line 7 in Alg. 1 returns true for label $l^{\prime}$, procedure SolveMHPP needs to be called for $l^{\prime}$ and SolveMHPP may return a new joint sequence $\gamma^{\prime}$. In this new sequence $\gamma^{\prime}$, it is possible that agents within $I_{C}(l)$ visits targets that are previously assigned to agents that are not in $I_{C}(l)$. However, DMS* does not let agents outside $I_{C}(l)$ to take all possible actions as defined in Equation (1). As a result, an optimal solution joint path may lie outside $\mathcal{G}^{W}_{sub}$ and DMS* thus loses the solution optimality guarantee when Assumption 2 is violated.
To ensure solution optimality when Assumption 2 is violated, same as in [8], one possible way is to back propagate the entire index set $I=\{1,2,\cdots,N\}$ as the conflict set when calling BackProp. This ensures that DMS* considers all possible actions for all agents when conflicts between agents are detected. However, in practice, this is computationally burdensome when $N$ is large, and limits the scalability of the approach. We therefore omit it from Alg. 1 for clarity.
IV-D3 Solution Bounded Sub-Optimality
A joint sequence $\gamma$ is $\epsilon$-bounded sub-optimal ($\epsilon\geq 0$) if $\max_{i\in I}cost(\gamma^{i})\leq(1+\epsilon)\max_{i\in I}cost(\gamma_{*}^{i})$, where $\gamma_{*}$ is an optimal joint sequence. Similarly, a solution joint path $\pi$ is $\epsilon$-bounded sub-optimal if its objective value $g\leq(1+\epsilon)g_{*}$, where $g_{*}$ is the true optimum of the given instance.
Theorem 3.
For a solvable instance, if (i) SolveMHPP returns a $\epsilon$-bounded sub-optimal joint sequence for any given label $l$ and (ii) Assumption 2 holds, then DMS* returns a $\epsilon$-bounded solution joint path.
Bounded sub-optimal joint sequences lead to inflated heuristic values for labels in DMS*. As DMS* select labels based on $f$-values as A* does, inflated heuristic values lead to bounded sub-optimal solutions for the problem [31].
In addition to using bounded sub-optimal joint sequences, DMS* is also able to use the conventional heuristic inflation technique [31] and provide bounded sub-optimal solution.
Theorem 4.
For a solvable instance, if Line 24 in Alg. 1 is replaced with $\vec{f}\leftarrow\vec{g}+(1+w)\vec{h}$ and Assumption 1 and 2 hold, then DMS* returns a $w$-bounded solution joint path.
V Experimental Results
We implement DMS* in Python, and use Google OR-Tool [^5] https://developers.google.com/optimization/routing/routing_tasks as the mHPP solver, which implements the SolveMHPP procedure that is required by DMS*. Limited by our knowledge on Google OR-Tool, we only consider the following type of assignment constraints, where any agent can visit all targets and goals. We use grid maps from the MAPF benchmark data set [32] and set up a 60-second runtime limit for each test instance, where each test instance contains the starts, targets and goals, which are all from the MAPF benchmark data set [32]. All tests run on a MacBook Pro with a Apple M2 Pro CPU and 16GM RAM. It is known from [10] that, for M*-based algorithms, a slightly inflated heuristic can often help improve the scalability of the approach with respect to the number of agents and we set $w=1.1$ in our tests. We set the number of targets $M=20,40,60,80$ in our tests. We compare DMS*, which includes the two proposed technique to defer target sequencing, against MS*, which is the baseline and does not have these two techniques. In other words, MS* here is a naive adaption of the existing MS* [8] algorithm to solve MCPF-max, by using the aforementioned Google OR-Tool to solve min-max mHPP for target sequencing.
Figure 3: The success rates and runtime of both DMS* and MS* (baseline) with varying number of agents and targets in a random 32x32 map. DMS* achieves higher success rates than MS* while requiring less runtime on average.
V-A Varying Number of Agents
We first fix the map to be Random 32x32, and vary the number of agents $N=5,10,15,20$. We measure the success rates, the average runtime to solution and the average runtime for target sequencing per instance. The averages are taken over all instances, including both succeeded instances and instances where the algorithm times out. As shown in Fig. 3, DMS* achieves higher success rates and lower runtime than the baseline MS* for all $N$s. As $M$ increases from $20$ to $80$, both algorithms require more runtime for target sequencing. As $N$ increases from $5$ to $20$, agents have higher density and are more likely to run into conflict with each other. As a result, both algorithms time out for more and more instances. Additionally, MS* spends almost all of its runtime in target sequencing, which indicates the computational burden caused by the frequent call of mHPP solver. In contrast, for DMS*, when $N=5,10$, DMS* spends most of the runtime in target sequencing, while as $N$ increases to $15,20$, DMS* spends more runtime in path planning.
Figure 4: The success rates and runtime of both DMS* and MS* (baseline) with varying number of targets in maps of different sizes. DMS* achieves higher success rates than MS* while requiring less runtime on average.
V-B Different Maps
We then fixed the number of agents to $N=10$ and test both algorithms in several maps of different sizes, as shown in Fig. 4. DMS* outperforms MS* in terms of success rates due to the alleviated computational burden for target sequencing. Furthermore, larger maps do not lead to lower success rates since larger maps can reduce the density of the agents and make the agents less likely to run into conflicts with each other. The Room 32x32 map is relatively more challenging than the other two maps since there are many narrow corridors which often lead to conflicts between the agents.
V-C Experiments with Mobile Robots
We also performed experiments with Husarion ROSbot 2666https://husarion.com/manuals/rosbot/, which are differential drive robots equipped with an ASUS Tinker Board and use a 2D lidar for localization. We used two such robots to perform tests in two different environments, as shown in Fig. 5 (a) and (b) having 2 and 3 targets, respectively. This test verifies that the paths planned by DMS* are executable on real robots. When the robots have large motion disturbance, such as delay or deviation from the planned path, additional techniques (such as [33]) will be needed to ensure collision-free execution of the paths, which is another research topic and is not the focus of this work. In our experiments, the robots run in slow speed and the motion disturbance of the robots are relatively small.
Figure 5: Test settings for ROSbot experiments. (a) and (b) show the real-world setup. (c) and (d) show the rviz window corresponding to (a) and (b) respectively. Intermediate targets are highlighted using pink boxes and the goals using green.
We then compare the performance of DMS* (which minimizes the makespan) and MS* (which minimizes the sum of arrival times) on the mobile robots. Both planners were tested on the same map with two intermediate targets as shown in Fig. 6. We observe that DMS* tends to distribute the targets evenly to the agents and plan paths to minimize the makespan, while MS* finds a solution wherein one robot visits all targets, and the other robot directly goes to its assigned destination. MS* seeks to minimize the sum of arrival times, which can lead to a large makespan.
Figure 6: Comparison of DMS* and MS* on mobile robots. (a) shows the test setting. (b) and (c) show the joint path planned by DMS* and MS* respectively. Both planners return different paths since they optimize different objective functions.
VI Conclusion and Future Work
This paper investigates a min-max variant of Multi-Agent Combinatorial Path Finding problem and develops DMS* algorithm to solve this problem. We analyze the properties of DMS* and test DMS* with up to 20 agents and 80 targets. We also conduct robot experiments to showcase the usage of DMS* on differential-drive robots.
For future work, one can investigate how to incorporate the existing conflict resolution techniques for MAPF into the DMS* to further improve its runtime efficiency. We also note from our robot experiments that, the uncertainty and disturbance in robot motion may affect the execution of the planned path, and one can develop fast online replanning version of DMS* to address the uncertainty and disturbance.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. 2120219 and 2120529. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.